# Reversibility in the Seifert-fibered spaces

### 03 Mar 2024 v1

@@ 1. Introduction
An element $a$ in a group $\Gamma$ is said to be ‘reversible’ if it is conjugate to its own inverse, and the conjugating elements are called ‘reverser’ or ‘reversing symmetry’. If the conjugating element is also an involution, then it is called ‘strongly reversible’ element. It has also been called by other terms like ‘reciprocals’ or ’real elements’ in other literature, e.g., [BR], [La] etc. In this paper our discussion is limited to some properties of reversible elements, and a discussion of reversible elements in some discrete geometric groups. Reversibility has its own geometric motivations and has been discussed widely in literature. For instance, in a Fuchsian group, the collection of conjugacy classes of reversible elements have a 1-1 correspondence to the conjugacy classes of infinite dihedral groups contained in that Fuchsian group. But in general, supposing $G$ is a 3-manifold group (the fundamental group of some 3-manifold), if $G$ is torsion free then a reversible element $g$ in $G$ is represented as a loop which is homotopic to its own inverse (i.e., it will form an embedded Klein bottle with the reversing symmetry). So, it can be considered that finding a reversible element with a reverser in a 3-manifold is equivalent to finding an embedded Klein bottle upto homotopy inside the manifold. In this paper we are specially interested in Seifert fibered manifolds since the fundamental group of a Seifert fibered manifold is a universal extension of Fuchsian groups or non-orientable 2-orbifold groups.
@@
$\pi_{1}(M)=$ $\langle a_{i},b_{i},c_{i},d_{i},h|a_{i}ha_{i}^{-1}=h^{\phi(a_{i})},$
$b_{i}hb_{i}^{-1}=h^{\phi(b_{i})},c_{i}hc_{i}^{-1}=h,d_{i}hd_{i}^{-1}=h^{\phi(d_{i})},$
$c_{i}^{\mu_{i}}=h^{\beta}_{i},\Pi[a_{i},b_{i}]\Pi c_{i}\Pi d_{i}h^{b}=1\rangle.$
@@
$\pi_{1}(M)=$ $\langle x_{i},c_{i},d_{i},h|x_{i}hx_{i}^{-1}=h^{\phi(x_{i})},$
$c_{i}hc_{i}^{-1}=h,d_{i}hd_{i}^{-1}=h^{\phi(d_{i})},$
$c_{i}^{\mu_{i}}=h^{\beta_{i}},\Pi x_{i}^{2}\Pi c_{i}\Pi d_{i}h^{b}=1\rangle.$
@@
From the representation of $\pi_{1}(M)$, we can see that the reversibility of the element $h$ is dependent on the classifying homomorphism $\phi$. If there is any generator $z$ of the form $a_{i},b_{i},x_{i},$ or $d_{i}$ such that $\phi(z)=-1$, then the relations in the definition of $\pi_{1}(M)$ give that $h$ is reversible, i.e., $zhz^{-1}=h^{\phi(z)}=h^{-1}$. Our aim is to classify all the reversible elements of $\pi_{1}(M)$. In Section 2, we discuss about the reversible elements of fundamental groups of two-dimensional hyperbolic orbifolds, and we classify the reversible elements of such groups. This classification is then applied to the fundamental groups of any Seifert-fibered space in Section 3. Section 4 then discusses an application of the results of Section 3 on the fundamental group of $B_{3}$, the braid group on $3$ strands. The possibility of using the same methods for any $B_{n}/\langle h\rangle$ is also discussed there.
@@ Acknowledgment
A. Das acknowledges full support from an NBHM research fellowship during this work. D. Das acknowledges full support from a UGC-NFSC research fellowship during this work. The authors thank Prof. Krishnendu Gongopadhyay and Dr. Soma Maity for helpful feedback.
@@ 2.1. Reversible classes of Fuchsian group
Proof. $\Gamma$ is in $\rm PSL(2,\mathbb{R})$, and so we will conjugate $\Gamma$ by a möbious map $f$ such that in the new group after conjugation, $h$ will be a hyperbolic element in the Poincaré disk where the hyperbolic axis of $h$ passes through $(0,0)$. Now, the reciprocator for $h$ is $g$, which is an involution having fixed point $z$. So, let us suppose $z$ does not lie on $I_{h}$ and $I_{h}$ is a straight line passing through the origin. Then, $z$ lies one side of the axis. In our earlier work [DG1, Lemma 2.2] we have shown that in the Fuchsian group, the reciprocators of a reciprocal element fix the hyperbolic axis of that particular hyperbolic element and interchange the fixed points of hyperbolic axis. Now, we apply the map $g$ to the disk. This map is a $\pi$-rotation along $I_{h}$ of the disk, and therefore the point $z$ must change its position. This is contradiction. So, the fixed point of $g$ must lie on $I_{h}$. ∎
@@ 3. Seifert fibered space
If $M$ has an orientable orbit surface, the only reversible elements in $\pi_{1}(M)$ can be the conjugates of the powers of $h$ and the conjugates of the elements of the form $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}$ where $\mu_{i}$ and $\mu_{j}$ are even, and $k$ is any element in $\pi_{1}(M)$, $\beta_{i}=\beta_{j}$ and $\phi(k)=-1$.
@@
$c^{\frac{\mu_{i}}{2}}c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}c^{-\frac{\mu_{i}}{2}}c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}$
$=c^{\mu_{i}}c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}$
$=h^{\beta_{i}}kh^{\beta_{j}}k^{-1}.$
Then, $h^{\beta_{i}}kh^{\beta_{j}}k^{-1}$ will be identity only if $\beta_{i}=\beta_{j}$ and $\phi(k)=-1.$ It will be same if we choose any other reverser for that particular reversible element since all the reversers are in the form of $x^{l}c^{\frac{\mu_{i}}{2}}.$ If $x$ is not reversible in $\pi_{1}(M)/\langle h\rangle$, a reverser of $x$ in $\pi_{1}(M)$, say $\rho$, must be a power of $h$, since the projection to $\pi_{1}(M)/\langle h\rangle$ must be constant. Then in $\pi_{1}(M)/\langle h\rangle$, $[x]=[x^{-1}]$. This implies that $[x^{2}]=1$. Then, $x^{2}$ is a power of $h$, and $x^{2}$ is also reversible with same reverser $\rho$ that $x$ is. Then,
@@
The previous lemma shows that there are no reversible elements in $\pi_{1}(M)$ other than those conjugate to $h$. As shown earlier, the reversibility of $h$ and its conjugates depends on the classifying homomorphism. This is summarised in the following statement:
Theorem 3.6.
If $M$ has an orientable orbit surface and $\phi$ is nontrivial, the reversible elements of $\pi_{1}(M)$ are the elements conjugate to $h$ and $c_{i}^{\mu_{i}}$ for all $i$ and the conjugates of the elements of the form $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}$ where $\mu_{i}$ and $\mu_{j}$ are even, and $k$ is any element in $\pi_{1}(M)$, with $\beta_{i}=\beta_{j}$ and $\phi(k)=-1$. $\pi_{1}(M)$ has no reversible elements if $\phi$ is trivial.
@@
Theorem 3.7.
If $M$ has an non-orientable orbit surface other than the Klein bottle or $\mathbb{RP}^{2}$, the reversible elements of $\pi_{1}(M)$ are either conjugate to any power of $h$ or $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}$ if $\mu_{i}$ and $\mu_{j}$ are even and $\phi(k)=-1$.
@@
Theorem 3.8.
The classification of the reversible elements in a Seifert-fibered space depends only on the classifying homomorphism.
@@ 4. Application to universal extension of Fuchsian groups
Here, $\pi_{1}(M)$ is the fundamental group of $M$, and $F$ is a Fuchsian group or discrete isometry group of $\rm PGL(2,\mathbb{R})$. So, $\pi_{1}(M)$ is then a universal extension of $F$, which can now be seen as $\pi_{1}(M)/\langle h\rangle$. This gives us a new view-point to consider certain groups as fundamental groups of a Seifert-fibered space. Let us consider $F=\rm PSL(2,\mathbb{Z})$ which is presented by $\langle a,b|a^{2},b^{3}\rangle.$ We have already mentioned the presentation of $\pi_{1}(M)$ for a given $M$ in 1 and 1. We can choose an orientable connected $3$-manifold $M$ in such a way that the $\pi_{1}(M)\cong B_{3}$. $B_{3}$ has a presentation as $\langle\sigma_{1}\sigma_{2}|\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2}\rangle$. The normal subgroup of $B_{3}$ is $\mathbb{Z}$ and it is generated by $(\sigma_{1}\sigma_{2})^{3}$. Thus, $B_{3}/\langle\sigma_{1}\sigma_{2}\rangle=\langle\sigma_{1}\sigma_{2}\sigma_{1},\sigma_{1}\sigma_{2}|(\sigma_{1}\sigma_{2})^{3},(\sigma_{1}\sigma_{2}\sigma_{1})^{2}\rangle=\rm PSL(2,\mathbb{Z})\cong\Delta(2,3,\infty)$. Choosing the $c_{i}$ and $\beta_{i}$ suitably, we can consider the Seifert-fibered manifold $M$ determined by $(O,o,0|1,(2,1),(3,1))$. The presentation of the fundamental group of $M$ is then
@@
$B_{3}$ has no reversible elements.
@@
Remark 4.2.
By the same argument, we can obtain that there is no reversible element in an universal extension of any triangle group.
Also note that $\rm SL(2,\mathbb{Z})=\langle\sigma_{1},\sigma_{2}|\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2},(\sigma_{1}\sigma_{2})^{6}\rangle\cong Mod(S_{1,1}).$
Remark 4.3.
The previous discussion concludes that the reversible elements of hyperbolic groups will not carry to the universal central extension in general, since triangle groups are hyperbolic groups but its universal extensions are not necessarily so.
@@
$B_{n}=\langle x,y|x^{n}=(xy)^{n-1},x^{i}yx^{-i}y=yx^{i}yx^{-i}\hbox{ where }i=2,3,\dots,m\rangle$
@@
From the above theorem, we obtain that the same idea that we have used for Seifert-fibered spaces will not work for $B_{n}$. So, we need to follow some different process.

### 06 Mar 2024 v2

@@ 1. Introduction
An element $a$ in a group $\Gamma$ is said to be ‘reversible’ if it is conjugate to its own inverse, and the conjugating elements are called ‘reverser’ or ‘reversing symmetry’. If the conjugating element is also an involution, then it is called ‘strongly reversible’ element. It has also been called by other terms like ‘reciprocals’ or ‘real elements’ in other literature, e.g., [BR], [La] etc. In this paper our discussion is limited to some properties of reversible elements, and a discussion of reversible elements in some discrete geometric groups. Reversibility has its own geometric motivations and has been discussed widely in literature. For instance, in a Fuchsian group, the collection of conjugacy classes of reversible elements have a 1-1 correspondence to the conjugacy classes of infinite dihedral groups contained in that Fuchsian group. But in general, supposing $G$ is a 3-manifold group (the fundamental group of some 3-manifold), if $G$ is torsion free then a reversible element $g$ in $G$ is represented as a loop which is homotopic to its own inverse (i.e., it will form an embedded Klein bottle with the reversing symmetry). So, it can be considered that finding a reversible element with a reverser in a 3-manifold is equivalent to finding an embedded Klein bottle upto homotopy inside the manifold. In this paper we are specially interested in Seifert fibered manifolds since the fundamental group of a Seifert fibered manifold is a universal extension of Fuchsian groups or non-orientable 2-orbifold groups.
@@
(1.1) $\pi_{1}(M)=$ $\langle a_{i},b_{i},c_{i},d_{i},h|a_{i}ha_{i}^{-1}=h^{\phi(a_{i})},$
(1.2) $b_{i}hb_{i}^{-1}=h^{\phi(b_{i})},c_{i}hc_{i}^{-1}=h,d_{i}hd_{i}^{-1}=h^{\phi(d_{i})},$
(1.3) $c_{i}^{\mu_{i}}=h^{\beta}_{i},\Pi[a_{i},b_{i}]\Pi c_{i}\Pi d_{i}h^{b}=1\rangle.$
@@
(1.4) $\pi_{1}(M)=$ $\langle x_{i},c_{i},d_{i},h|x_{i}hx_{i}^{-1}=h^{\phi(x_{i})},c_{i}hc_{i}^{-1}=h,d_{i}hd_{i}^{-1}=h^{\phi(d_{i})},$
(1.5) $c_{i}^{\mu_{i}}=h^{\beta_{i}},\Pi x_{i}^{2}\Pi c_{i}\Pi d_{i}h^{b}=1\rangle.$
@@
From the representation of $\pi_{1}(M)$, we can see that the reversibility of the element $h$ is dependent on the classifying homomorphism $\phi$. If there is any generator $z$ of the form $a_{i},b_{i},x_{i},$ or $d_{i}$ such that $\phi(z)=-1$, then the relations in the definition of $\pi_{1}(M)$ give that $h$ is reversible, i.e., $zhz^{-1}=h^{\phi(z)}=h^{-1}$. Our aim is to classify all the reversible elements of $\pi_{1}(M)$. The main result of our paper is:
Theorem 1.1.
If $M$ has an orientable orbit surface and $\phi$ is nontrivial, the reversible elements of $\pi_{1}(M)$ are the elements conjugate to $h$ and $c_{i}^{\mu_{i}}$ for all $i$ and the conjugates of the elements of either of the form $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}$ with $\phi(k)=-1$ or of the form $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{-\mu_{j}}{2}}k^{-1}$ with $\phi(k)=1$, where $\mu_{i}$ and $\mu_{j}$ are even, and $k$ is any element in $\pi_{1}(M)$, $\beta_{i}=\beta_{j}$. $\pi_{1}(M)$ has reversible elements as the elements conjugate to $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{-\mu_{j}}{2}}k^{-1}$ if $\phi$ is trivial.
We also similarly classify the reversible elements in case of Seifert fibered spaces with non-orientable orbit surfaces. In Section 2, we discuss about the reversible elements of fundamental groups of two-dimensional hyperbolic orbifolds, and we classify the reversible elements of such groups. This classification is then applied to the fundamental groups of any Seifert-fibered space in Section 3. Section 4 then discusses an application of the results of Section 3 on the fundamental group of $B_{3}$, the braid group on $3$ strands. The possibility of using the same methods for any $B_{n}/\langle h\rangle$ is also discussed there. After completion of the first draft, the authors were informed by Prof. Masakazu Teragaito regarding the article [HMT] written by Himeno, Motegi, and Teragaito, where they classify 3-manifolds whose fundamental groups admit generalised torsion elements of order two, and further classify such elements. Their proof is topological and completely different from the proof in this paper. The two articles were written completely independently and at around the same time. The proof in [HMT] crucially uses important results by Jaco-Shalen and Hass, while the techniques in this article are simpler by comparison.
@@ Acknowledgment
The authors would like to acknowledge and thank Prof. Masakazu Teragaito for his helpful comments and for pointing out an error in the first draft. The authors further thank Prof. Krishnendu Gongopadhyay and Dr. Soma Maity for helpful feedback. A. Das acknowledges full support from an NBHM research fellowship during this work. D. Das acknowledges full support from a UGC-NFSC research fellowship during this work.
@@ 2.1. Reversible classes of Fuchsian group
Proof. $\Gamma$ is in $\rm PSL(2,\mathbb{R})$, and so we will conjugate $\Gamma$ by a Möbius map $f$ such that in the new group after conjugation, $h$ will be a hyperbolic element in the Poincaré disk where the hyperbolic axis of $h$ passes through $(0,0)$. Now, the reciprocator for $h$ is $g$, which is an involution having fixed point $z$. So, let us suppose $z$ does not lie on $I_{h}$ and $I_{h}$ is a straight line passing through the origin. Then, $z$ lies one side of the axis. In our earlier work [DG1, Lemma 2.2] we have shown that in the Fuchsian group, the reciprocators of a reciprocal element fix the hyperbolic axis of that particular hyperbolic element and interchange the fixed points of hyperbolic axis. Now, we apply the map $g$ to the disk. This map is a $\pi$-rotation along $I_{h}$ of the disk, and therefore the point $z$ must change its position. This is contradiction. So, the fixed point of $g$ must lie on $I_{h}$. ∎
@@ 3. Seifert fibered space
If $M$ has an orientable orbit surface, the only reversible elements in $\pi_{1}(M)$ can be the conjugates of the powers of $h$ and the conjugates of the elements either of the form $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}$ with $\phi(k)=-1$ or of the form $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{-\mu_{j}}{2}}k^{-1}$ with $\phi(k)=1$, where $\mu_{i}$ and $\mu_{j}$ are even, and $k$ is any element in $\pi_{1}(M)$, $\beta_{i}=\beta_{j}$.
@@
$c_{i}^{\frac{\mu_{i}}{2}}c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}c_{i}^{-\frac{\mu_{i}}{2}}c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}=c_{i}^{\mu_{i}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}=h^{\beta_{i}}kh^{\beta_{j}}k^{-1}.$
Then, $h^{\beta_{i}}kh^{\beta_{j}}k^{-1}$ will be identity only if $\beta_{i}=\beta_{j}$ and $\phi(k)=-1.$ It will be same if we choose any other reverser for that particular reversible element since all the reversers are in the form of $x^{l}c^{\frac{\mu_{i}}{2}}.$ In the similar way, we can show that $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{-\frac{\mu_{j}}{2}}k^{-1}$ is reversible if $\phi(k)=1$. If $x$ is not reversible in $\pi_{1}(M)/\langle h\rangle$, a reverser of $x$ in $\pi_{1}(M)$, say $\rho$, must be a power of $h$, since the projection to $\pi_{1}(M)/\langle h\rangle$ must be constant. Then in $\pi_{1}(M)/\langle h\rangle$, $[x]=[x^{-1}]$. This implies that $[x^{2}]=1$. Then, $x^{2}$ is a power of $h$, and $x^{2}$ is also reversible with same reverser $\rho$ that $x$ is. Then,
@@
3.1. Proof of Theorem 1.1
The previous lemma shows that there are no reversible elements in $\pi_{1}(M)$ other than those conjugate to $h$. As shown earlier, the reversibility of $h$ and its conjugates depends on the classifying homomorphism. This is summarised in the statement of Theorem 1.1. ∎
@@
Corollary 3.6.
If $M$ has an non-orientable orbit surface other than the Klein bottle or $\mathbb{RP}^{2}$, the reversible elements of $\pi_{1}(M)$ are either conjugate to any power of $h$ or either of the form $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{\mu_{j}}{2}}k^{-1}$ with $\phi(k)=-1$ or of the form $c_{i}^{\frac{\mu_{i}}{2}}kc_{j}^{\frac{-\mu_{j}}{2}}k^{-1}$ with $\phi(k)=1$, where $\mu_{i}$ and $\mu_{j}$ are even, and $k$ is any element in $\pi_{1}(M)$, $\beta_{i}=\beta_{j}$.
@@
Remark 3.7.
The classification of the reversible elements in a given Seifert-fibered space depends only on the classifying homomorphism.
@@ 4. Application to universal extension of Fuchsian groups
Here, $\pi_{1}(M)$ is the fundamental group of $M$, and $F$ is a Fuchsian group or discrete isometry group of $\rm PGL(2,\mathbb{R})$. So, $\pi_{1}(M)$ is then a universal extension of $F$, which can now be seen as $\pi_{1}(M)/\langle h\rangle$. This gives us a new view-point to consider certain groups as fundamental groups of a Seifert-fibered space. Let us consider $F=\rm PSL(2,\mathbb{Z})$ which is presented by $\langle a,b|a^{2},b^{3}\rangle.$ We have already mentioned the presentation of $\pi_{1}(M)$ for a given $M$ in Equation (1) and Equation (1). We can choose an orientable connected $3$-manifold $M$ in such a way that the $\pi_{1}(M)\cong B_{3}$. $B_{3}$ has a presentation as $\langle\sigma_{1}\sigma_{2}|\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2}\rangle$. The normal subgroup of $B_{3}$ is $\mathbb{Z}$ and it is generated by $(\sigma_{1}\sigma_{2})^{3}$. Thus, $B_{3}/\langle\sigma_{1}\sigma_{2}\rangle=\langle\sigma_{1}\sigma_{2}\sigma_{1},\sigma_{1}\sigma_{2}|(\sigma_{1}\sigma_{2})^{3},(\sigma_{1}\sigma_{2}\sigma_{1})^{2}\rangle=\rm PSL(2,\mathbb{Z})\cong\Delta(2,3,\infty)$. Choosing the $c_{i}$ and $\beta_{i}$ suitably, we can consider the Seifert-fibered manifold $M$ determined by $(O,o,0|1,(2,1),(3,1))$. The presentation of the fundamental group of $M$ is then
@@
The only reversible elements of $B_{3}$ are conjugates to $[\sigma_{1}\sigma_{2}\sigma_{1},k]$ for any element $k$ in $B_{3}$.
@@
(4.4) $B_{n}=\langle x,y|x^{n}=(xy)^{n-1},x^{i}yx^{-i}y=yx^{i}yx^{-i}\hbox{ where }i=2,3,\dots,m\rangle$
@@
However, there always exits some reversible elements in $B_{n}$ by the following remark. Here, we are assuming the presentation of $B_{n}$ as given in (4.4).
@@
From Corollary 4.1, there are some reversible elements in $B_{n}$ which are conjugate to the elements of the form $[a,b]$, where the image of $a$ is an involution in the quotient space $B_{n}/\langle x^{n}\rangle$ and $b$ is any other element in $B_{n}$. This is a sufficient condition to be a reversible element in $B_{n}$.
Remark 4.4.
From the above theorem, we obtain that the same idea that we have used for Seifert-fibered spaces will not work for $B_{n}$. Furthermore, by [HMT, Theorem 1.3] and Theorem 4.2, $B_{n}$ cannot be a fundamental group of any $3$-manifold. This has been proven earlier using different techniques, for example in [Go] where they use the theory of coherent groups, however this gives a different perspective using reversible elements.